JAVA question: Suppose we are performing a binary search on a sorted array called numbers initialized as follows:

// index           0   1   2   3   4   5   6   7   8   9  10  11  12  13
int[] numbers = {-30, -9, -6, -4, -2, -1,  0,  2,  4, 10, 12, 17, 22, 30};
// search for the value -5
int index = binarySearch(numbers, -5);

Write the indexes of the elements that would be examined by the binary search (the mid values in our algorithm's code) and write the value that would be returned from the search. Assume that we are using the binary search algorithm shown in chapter 13 of the Building Java Programs 4th Edition textbook.

Indexes examined:

6, 2, 4, 3

value returned:

-1

This is the answer, but can someone please explain why all of those indexes are examined? Why not just indexes 2 and 3?
Answer #1

First it will find the mid element as

left =0 and right =13 so 0+13-1/2= 6 so 6 is examined

now numbers[6]>num

so now we will go left and mid-1

left=0 and right=5

mid= 0-5-1/2=2

so arr[2]<-5 so

so now we will go mid+1 and right-1

so left=3 and right 4

so 3+4+1/2=4

so similar to others also

Note : Please comment below if you have concerns. I am here to help you

If you like my answer please rate and help me it is very Imp for me